Operation Summary for November 1, 2010 - November 3, 2010
Monday Nov. 1
The day shift began with Operations monitoring the Tevatron (TeV) store 8227 (79.05E30) and the Recycler stash (147.4E10), with the Antiproton Source (Pbar) stacking (196.9E10), with the Main Injector (MI), Linac, and Booster operating normally, with NuMI off for target work, with the Pelletron off for repairs, with the MiniBooNE and Meson FTBF CALICE T-978 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.
During the day shift, Operations terminated store 8227 at 9:43 AM, put the TeV into HEP shot setup, and established store 8228 at 10:55 with an initial luminosity (AIL) of 129E30. Operations established beam to NuMI at 11:01 AM. Unfortunately, NuMI went down for two hours due to a high target LCW conductivity alarm. Operations reported that a loose cable connection in the low level RF system caused some Booster downtime. Pelletron repairs continued throughout the shift.
During the evening shift, Operations disabled beam to T-978 at 6 PM. TeV personnel began a crystal collimator end-of-store study at 6:30 PM, which lasted three hours. Operations terminated store 8228 at 10:13 PM, put the TeV into shot setup, and established store 8229 at 11:36 PM with an AIL of 131.2E30. Repairs on the Pelletron continued.
Tuesday Nov. 2
The midnight shift began with Operations monitoring store 8229 (112.81E30) and the stash (146.9E10), with the Pelletron off for repairs, with Pbar stacking (53.4E10), and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations established beam to T-978 at 4 AM. Electron Cooling (ECool) personnel completed their Pelletron repairs and Mechanical technicians had the insulating gas (SF6) transferred back into the Pelletron by 6:50 AM. Operators searched and secured MI-31 and ECool personnel began the Pelletron turn on process.
During the day shift, Operations terminated store 8229 at 8:18 AM and kicker personnel began work on the Tevatron abort kickers. Operations reported that the Pelletron resumed normal operations at 8:52 AM and Pbar resumed antiproton transfers to the Recycler. Part of the D0 collision hall systems (emergency generator, cryo system power, chilled water, solenoid, and torroid) lost power at 9:40 AM; operators prepared the enclosure for access, but D0 personnel resolved the problem without having to go in. The Tevatron quenched at 12:29 PM during the D0 low beta magnet turn on; recovery continued to the end of shift. TeV kicker personnel completed their work on the TeV abort kicker at 12:30 PM. Operations reported minor downtimes due to Linac and MI RF system trips.
During the evening shift, Operations completed the Tevatron turn on and checkout at 5:15 PM, put the TeV into shot setup at 5:23 PM, and established store 8231 at 6:45 PM with an AIL of 317.1E30. Operations halted stacking for 1.5 hours due to a dipole magnet power supply tripping off on a phase unbalance; Operations called in power supply personnel. Toward the end of the shift, the Booster Critical Device began showing intermittent digital status via ACNET; investigations lasted until the end of shift.
Wednesday Nov. 3
The midnight shift began with Operations monitoring store 8231 (137.36E30) and the stash (191.4E10), with Pbar stacking (12.2E10), and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations terminated store 8231 at 7:34 AM. Operators prepared D0 for a one-hour access. Operators tuned on the Booster and MI machines throughout most of the shift.
The Plans for Wednesday
The plans for today are to stack, stash, and store.
Comments and Suggestions
|last modified 11/03/2010 email Fermilab|
|Security, Privacy, Legal|